Moment of inertia of a hollow sphere of mass M and diameter D about its diameter 

A

MD2/5

B

2MD2/5

C

MD2/6

D

5MD2/2

 

Grade:12th pass

1 Answers

Adarsh Pal
3 years ago

 

Moment of inertia of a hollow sphere of mass M and diameter D about its diameter is 2MR2/3
Since, R = D/2, putting the value of R in the above formula,
I = 2 M D2/ 4 x 3 = MD2/6
So, the correct answer is 'MD2/6' i.e C option